# Lattice tuning of charge and spin transport in $\beta_{12}$-borophene nanoribbons

### v1

@@ IV Conclusion
We have systematically investigated how electron–phonon coupling (EPC) influences the electronic transport properties of $\beta_{12}$-borophene nanoribbons (BNRs). Using an ab initio-based tight-binding model combined with the Holstein approximation and the nonequilibrium Green’s function–Landauer formalism, we demonstrate that EPC significantly reshapes both the magnitude and the energy dependence of conduction pathways. Zigzag-edged BNRs exhibit spin polarization, whereas other edge configurations such as line–line remain nonmagnetic and display higher conductance. EPC consistently enhances the overall current through phonon-assisted inelastic processes, redistributes spectral weights in the density of states, and broadens differential conductance features. These effects underline the importance of lattice dynamics and edge-specific electronic structures, positioning EPC and edge engineering as critical strategies for tuning electronic and spintronic performance in borophene-based nanoscale devices.
@@ V Acknowledgments
M.D. gratefully acknowledges Soheil Ershadrad for valuable discussions and support with Wannier90 calculations. B.S. acknowledges financial support from Swedish Research Council (grant no. 2022-04309), STINT Mobility Grant for Internationalization (grant no. MG2022-9386) and DST-SPARC, India (Ref. No. SPARC/2019-2020/P1879/SL). The computations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) at UPPMAX (NAISS 2024/5-258) and at NSC and PDC (NAISS 2023/3-42) partially funded by the Swedish Research Council through grant agreement no. 2022-06725. B.S. also acknowledges the allocation of supercomputing hours granted by the EuroHPC JU Development Access call in LUMI-C supercomputer (grant no. EHPC-DEV-2024D04-071) in Finland.

### v2

@@ IV Conclusion
We have systematically investigated how EPC influences the spin and charge transport properties of $\beta_{12}$-BNRs. Using an ab initio-based tight-binding model combined with the Holstein approximation and the nonequilibrium Green’s function–Landauer formalism, we demonstrate that EPC significantly reshapes both the magnitude and the energy dependence of conduction pathways. Zigzag-edged BNRs enhances the weak inherent spin polarization, whereas other edge configurations such as line–line remain nonmagnetic, while they display higher conductance. EPC consistently enhances the overall current through phonon-assisted inelastic processes, redistributes spectral weights in the density of states, and broadens differential conductance features. These effects underline the importance of lattice degree of freedom and edge-specific electronic structures, positioning EPC and edge engineering as critical strategies for tuning electronic and spintronic performance in borophene-based nanoscale devices.
@@ V Acknowledgments
B.S. acknowledges financial support from Swedish Research Council (grant no. 2022-04309), STINT Mobility Grant for Internationalization (grant no. MG2022-9386) and DST-SPARC, India (Ref. No. SPARC/2019-2020/P1879/SL). The computations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) at UPPMAX (NAISS 2024/5-258) and at NSC and PDC (NAISS 2023/3-42) partially funded by the Swedish Research Council through grant agreement no. 2022-06725. B.S. also acknowledges the allocation of supercomputing hours granted by the EuroHPC JU Development Access call in LUMI-C supercomputer (grant no. EHPC-DEV-2024D04-071) in Finland.